40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 5939     Entry time: Fri Nov 18 01:27:04 2011
Author: Den 
Type: Update 
Category: IOO 
Subject: MC locked 

[Mirko, Den]

While the MC was unlocked (and the local damping off) we've measured the coherence between GUR1_X and OSEM sensors. It was rather high, close to 1 at frequencies 0.1 - 1 Hz. That means that stack does not kill all coherence between seismic noise and mirror motion.

Then we've turned on the local damping and measured the coherence again between GUR1_X and OSEM sensors. It decreased due to some noise and was on the level of ~0.5. We did reduced the motion between the mirror and the frame by local damping but it is not obvious that we lost some coherence due to this effect. Probably, actuator adds some noise.

When we locked the MC, we did not see any coherence at 0.1 - 1 Hz between GUR1_X or STS1_X and OSEM sensors of MC1 and MC3 but we did see with MC2. The MC1 sensor was fixed by Suresh.

 

Attachment 1: cohnolocalpumping-crop_4.pdf  34 kB  | Hide | Hide all
cohnolocalpumping-crop_4.pdf
Attachment 2: cohlocalpumping4-crop.pdf  34 kB  | Hide | Hide all
cohlocalpumping4-crop.pdf
Attachment 3: cohlock4-crop.pdf  34 kB  | Hide | Hide all
cohlock4-crop.pdf
ELOG V3.1.3-